# DeepFRC: An End-to-End Deep Learning Model for Functional Registration and Classification

### v2

@@ Abstract
Functional data - observations in the form of curves or trajectories - arise in diverse domains such as biomedical sensing, motion capture, and handwriting recognition. A core challenge in functional data analysis (FDA) is accounting for phase variability, where misaligned temporal patterns hinder accurate inference. We introduce DeepFRC, an end-to-end deep learning framework for joint functional registration and classification. Unlike conventional approaches that decouple alignment and prediction, DeepFRC integrates class-aware elastic warping and a learnable basis representation into a unified architecture. This design enables temporal alignment and dimensionality reduction to be jointly optimized with classification, improving both interpretability and accuracy. We establish the first theoretical connection between alignment quality and generalization error, and validate our model on synthetic and real-world benchmarks. DeepFRC consistently outperforms state-of-the-art methods, especially in scenarios with complex temporal misalignment. Code is available at: https://github.com/Drivergo-93589/DeepFRC.
@@ 1.3 Contributions
This work introduces a unified framework for functional registration and classification with the following key contributions: (1) Model Innovation: We propose (a) a learnable basis representation that preserves infinite-dimensional structure, (b) a class-aware registration loss tailored for multi-class alignment, and (c) a joint training strategy that enables mutual reinforcement between alignment and prediction - surpassing prior approaches (Lohit et al., 2019; Chen and Srivastava, 2021). (2) Theoretical Foundation: We establish the first theoretical connection between registration fidelity and classification generalization, offering new insights into how structured alignment impacts downstream learning. (3) Empirical Validation: Extensive experiments on both synthetic and real datasets demonstrate the model’s ability to jointly optimize registration and classification. DeepFRC consistently outperforms state-of-the-art baselines, with ablation studies validating the necessity of its integrated design. The paper follows the order of contributions outlined above.
@@ 2 The Model
Consider the data $\big{\{}(x_{i}(\bm{t}_{i}),y_{i})\big{\}}_{i=1}^{N}$, where $x_{i}(\bm{t}_{i})=\big{(}x_{i}(t_{i1}),\dots,x_{i}(t_{in})\big{)}^{\intercal}$ represents the observed values of functional data $x(t)$ at $n$ time points for the $i$-th sample, and $y_{i}\in\{1,\dots,C\}$ denotes its class label. The functional data $x(t)$ is assumed to be a mean-square continuous process defined over $t\in[0,1]$, and the observations are misaligned. We hypothesize the existence of an underlying nonlinear time warping map $\gamma:t\mapsto\tilde{t}$ and a prediction function $f:x(\tilde{t})\mapsto y$. Our goal is to estimate $\gamma$ and $f$ simultaneously by an end-to-end deep learning model. This allows us to algin new functional data $x^{*}(\bm{t}^{*})$ using $\gamma(\bm{t}^{*})$ and classify it with $y^{*}=f(x^{*}(\gamma(\bm{t}^{*})))$. In the following, we detail the development of our deep neural networks to approximate both mappings $\gamma$ and $f$.
2.1 1D CNN for Time Warping
Our approach firstly utilizes 1D convolutional neural networks (1D CNNs) to extract temporal features from functional data, capitalizing on their demonstrated effectiveness in capturing temporal patterns (Kiranyaz et al., 2016; Ince et al., 2016; Chen and Srivastava, 2021; Kiranyaz et al., 2021). These networks process observed functional data points to produce latent features, which are then used to construct time-warping functions, enabling better registration of functional data.
Given an input sequence $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n}$, we extract time-warping features $\tau(x_{i}(\bm{t}_{i}))\in\mathbb{R}^{n}$ using a multi-layer 1D CNN. Each layer applies a 1D convolution followed by a ReLU activation:
$\bm{h}^{(l)}=\text{ReLU}(W^{(l)}\circledast\bm{h}^{(l-1)}+B^{(l)}),\quad l=1,\ldots,l_{1}-1,$ (1)
where $\bm{h}^{(0)}=x_{i}(\bm{t}_{i})$. The final layer is a fully connected transformation:
$\tau(x_{i}(\bm{t}_{i}))=\text{ReLU}(W^{(l_{1})}\bm{h}^{(l_{1}-1)}+\bm{b}^{(l_{1})}).$ (2)
Each convolutional layer has learnable weights $W^{(l)}$ and biases $B^{(l)}$. Following Chen and Srivastava (2021), we set $l_{1}=4$ and use three hidden 1D‐CNN blocks (kernel size $3$; channels $16\!\rightarrow\!32\!\rightarrow\!64$). This encoder, parameterised by $\Theta_{1}$, outputs the warping field $\tau(x_{i}(\bm{t}_{i}))$ that captures temporal structure.
To construct a boundary-preserving diffeomorphic warping function, the latent features $\tau(x_{i}(\bm{t}_{i}))$ are transformed as:
$\tilde{\gamma}_{i}(t_{ij})=\frac{\sum_{\mu=0}^{j}\tau^{2}_{i\mu}}{\sum_{\nu=0}^{n}\tau^{2}_{i\nu}},\hskip 5.69054ptj=1,\dots,n,$ (3)
where $t_{i0}=0$, $t_{in}=1,\tau_{i0}=0$, and $\tau_{i\mu}(\mu>0)$ is the $\mu$-th element of $\tau(x_{i}(\bm{t}_{i}))$. This enures that $\tilde{\gamma}_{i}(0)=0,\tilde{\gamma}_{i}(1)=1$ and $\frac{d(\tilde{\gamma}_{i}(t))}{dt}>0$, satisfying the requirements of a valid warping function (Chen and Srivastava, 2021). To reduce roughness and preserve the geometric structure, the warping function is further smoothed:
@@
This smoothing step ensures a smooth and well-structured warping function (Chen and Srivastava, 2021).
2.2 Basis Representation for Aligned Function
By applying the time warping $\gamma_{i}(t_{ij})$ to the functional data $x_{i}(\bm{t}_{i})$, the aligned functional data $\tilde{x}_{i}(t)$ can be obtained using 1D linear interpolation on the points $\big{\{}\big{(}\gamma_{i}(t_{ij}),x_{i}(t_{ij})\big{)}\big{\}}_{j=1}^{n}$:
$\begin{split}\tilde{x}_{i}(t)\triangleq\sum_{r=0}^{n-1}\Big{[}x(t_{ir})+\frac{x_{i}(t_{i(r+1)})-x_{i}(t_{ir})}{\gamma_{i}(t_{i(r+1)})-\gamma_{i}(t_{ir})}(t-\gamma_{i}(t_{ir}))\Big{]}\cdot\mathds{1}_{\{\gamma_{i}(t_{ir})\leq t\leq\gamma_{i}(t_{i(r+1)})\}},\end{split}$ (5)
where $\mathds{1}_{\{\cdot\}}$ is the indicator function. We choose this linear interpolation because it is both stable and computationally efficient (Berrut and Trefethen, 2004).
While traditional methods often discretize $\tilde{x}_{i}(t)$ into a high-dimensional vector $\tilde{\bm{x}}_{i}$ for input into neural networks, this approach suffers from the curse of dimensionality and fails to capture the smoothness of functional data, especially in noisy settings (Yao et al., 2021). To address these issues, we approximate $\tilde{x}_{i}(t)$ using basis expansion:
@@ 2.2 Basis Representation for Aligned Function
where $\{\phi_{j}(t)\}_{j=1}^{K}$ are set to Fourier basis functions and $\{c_{ij}\}_{j=1}^{K}$ are coefficients to estimate. Using least squares, the coefficients $\bm{c}_{i}\triangleq(c_{i1},\dots,c_{iK})$ are estimated as:
[…]
2.3 Fully Connected Neural Network Classifier
To classify the $i$-th sample, we use a fully connected feedforward neural network that takes the input feature vector $\tilde{\bm{c}}_{i}\in\mathbb{R}^{K}$. The network has $l_{2}-1$ hidden layers with ReLU activations, followed by a softmax output layer:
$\bm{h}^{(l)}=\text{ReLU}(W^{(l)}\bm{h}^{(l-1)}+\bm{b}^{(l)}),\quad l=l_{1}+1,\dots,l_{1}+l_{2}-1.$ (8)
Here, $W^{(l)}$ and $\bm{b}^{(l)}$ are learnable weights and biases. The output $\psi(\tilde{\bm{c}}_{i})\in\mathbb{R}^{C}$ gives the class probabilities. The predicted label is:
$\psi(\tilde{\bm{c}}_{i})=\text{Softmax}(W^{(l_{1}+l_{2})}\bm{h}^{(l_{1}+l_{2}-1)}+\bm{b}^{(l_{1}+l_{2})}).$ (9)
The classifier is a three-layer MLP ($l_{2}=3$) with $8$ and $4$ hidden units, parameterised by $\Theta_{2}$.
2.4 Objective Function Formulation
We denote the overall parameters to estimate as $\Theta=\{\Theta_{1},\Theta_{2}\}$. Given training data $\{(x_{i}(\bm{t}_{i}),y_{i})\}_{i=1}^{N}$ (where $y_{i}\in\{1,\dots,C\}$), the objective function consists of two parts for estimating $\Theta$.
Class-aware Registration Loss.
We adopt the elastic Riemannian metric square-root velocity function (SRVF) (Srivastava and Klassen, 2016) for registration, which overcomes limitations of the $\mathds{L}^{2}$ metric such as lack of invariance to warping, asymmetry, and the pinching effect (Marron et al., 2015; Srivastava and Klassen, 2016). The SRVF of $x(t)$ is defined as $q(t)=\text{sign}(\dot{x}(t))\sqrt{|\dot{x}(t)|}$, and for the warped function $x(\gamma(t))$, it becomes $(q\star\gamma)(t)\triangleq q(\gamma(t))\sqrt{\dot{\gamma}(t)}$ (Joshi et al., 2007; Srivastava and Klassen, 2016). For the $i$-th sample, we denote the observed SRVF as $Q_{i}(\bm{\gamma}_{i})=\big{(}(q_{i}\star\gamma_{i})(t_{i1}),\dots,(q_{i}\star\gamma_{i})(t_{in})\big{)}^{\intercal}$. Inspired by the work of Chen and Srivastava (2021), we define the class-aware registration loss as:
$\begin{split}\mathcal{L}_{1}(\Theta_{1})={{\sum}}\limits_{j=1}^{C}\frac{\sum\limits_{i\in\{y_{i}=j\}}\lVert Q_{i}(\bm{\gamma}_{i})-\bar{Q}^{(j)}\rVert}{N^{(j)}}+\alpha\sum\limits_{1\leq u<v\leq C}\lVert\bar{Q}^{(u)}-\bar{Q}^{(v)}\rVert^{-1},\end{split}$ (10)
[…]
@@ 2.4 Objective Function Formulation
The cross-entropy loss is given by
$\mathcal{L}_{2}(\Theta)=-\frac{1}{N}\sum\limits_{i=1}^{N}\sum_{j=1}^{C}y_{ij}\text{log}\psi_{ij},$ (11)
where the label $y_{i}$ is represented as a one-hot encoded vector $(y_{i1},\dots,y_{iC})^{\intercal}$ where all elements are set to zero except for the element that corresponds to the active class, which is set to one.
The combined objective is:
@@
where $\beta$ balances registration and classification. Minimizing $\mathcal{L}_{1}$ aligns functions across classes while aiding classification, especially for similar classes, while minimizing $\mathcal{L}_{2}$ predicts class labels and improves alignment by enhancing class separation.
Integrating Subsections 2.1-2.4, we propose an end-to-end deep learning model, termed DeepFRC, for simultaneous functional registration and classification (Figure 1).
Figure 1: Overview of DeepFRC. The light blue region highlights the 1D convolutional module. The three blue arrows (left to right) denote: flattening of the $(l_{1}{-}1)$th layer matrix, the time warping (TW) operation, and the basis projection of the aligned function. “TW” and “DR” denote time warping and dimensionality reduction, respectively. We set $l_{1}=4,l_{2}=3$.
@@ 3 Theoretical Analysis
The phase variability is modeled using monotone and smooth transformations within the set $\Gamma=\{\gamma:[0,1]\to[0,1]\mid\gamma(0)=0,\gamma(1)=1,\dot{\gamma}>0\}$. Under the framework of elastic functional data analysis (EFDA) (Srivastava et al., 2011), the optimal warpings are defined as: $\bm{\gamma}^{*}=\underset{\{\gamma_{i}\}}{\operatorname*{argmin}}Q_{\text{reg}}(\bm{\gamma})\triangleq\underset{\{\gamma_{i}\}}{\operatorname*{argmin}}{{\sum}}_{j=1}^{C}\frac{\sum\limits_{i\in\{y_{i}=j\}}\lVert Q_{i}(\bm{\gamma}_{i})-\bar{Q}^{(j)}\rVert}{N^{(j)}},\gamma_{i}\in\Gamma$, where “reg” means “registration”. For any other estimated warping $\hat{\bm{\gamma}}$, the registration error can be quantified as $\Delta Q_{\text{reg}}(\bm{\gamma}^{*},\hat{\bm{\gamma}})\triangleq|Q_{\text{reg}}(\bm{\gamma}^{*})-Q_{\text{reg}}(\hat{\bm{\gamma}})|.$
@@
While Theorem 3.1 guarantees low registration error in theory, verifying this practically is challenging since the true warpings $\bm{\gamma}^{*}$ are not easily obtainable. This is due to the difficulty of achieving global optimization for $\min(Q_{\text{reg}}(\bm{\gamma}))$. However, the validity of the result can be assessed through simulation studies, where true warpings are provided.
@@
$\Delta R_{\text{gen}}(\hat{\Theta})=|\mathds{E}_{S,A}[R(\hat{\Theta})-R_{n}(\hat{\Theta})]|,$
@@
Assumption (i) in Theorem 3.3 is reasonable and easily satisfied. First, each class should exhibit a unique alignment pattern, meaning the corresponding $\bar{Q}$ values are distinct across classes. Second, by modifying the softmax function in the last layer of DeepFRC as: $\sigma(x_{j})=\frac{\text{exp}(x_{j})+\beta}{\sum_{i=1}^{C}(\text{exp}(x_{i})+\beta)}$, it is possible to select a constant $\beta$ such that each probability $\psi_{ij}$ is at least $\epsilon_{0}$.

### v3

@@ Abstract
Functional data, representing curves or trajectories, are ubiquitous in fields like biomedicine and motion analysis. A fundamental challenge is phase variability—temporal misalignments that obscure underlying patterns and degrade model performance. Current methods often address registration (alignment) and classification as separate, sequential tasks. This paper introduces DeepFRC, an end-to-end deep learning framework that jointly learns diffeomorphic warping functions and a classifier within a unified architecture. DeepFRC combines a neural deformation operator for elastic alignment, a spectral representation using Fourier basis for smooth functional embedding, and a class-aware contrastive loss that promotes both intra-class coherence and inter-class separation. We provide the first theoretical guarantees for such a joint model, proving its ability to approximate optimal warpings and establishing a data-dependent generalization bound that formally links registration fidelity to classification performance. Extensive experiments on synthetic and real-world datasets demonstrate that DeepFRC consistently outperforms state-of-the-art methods in both alignment quality and classification accuracy, while ablation studies validate the synergy of its components. DeepFRC also shows notable robustness to noise, missing data, and varying dataset scales. Code is available at https://github.com/Drivergo-93589/DeepFRC.
@@ 1.3 Contributions
Our main contributions are: (1) Unified Architecture for Joint Learning. We introduce the first end-to-end deep learning model that integrates a neural deformation operator for diffeomorphic warping, a spectral representation for smooth functional encoding, and a class-aware contrastive loss, enabling mutual reinforcement between alignment and prediction. (2) Theoretical Foundations. We establish the first theoretical guarantees for such a model, proving approximation capabilities and providing a generalization bound linking registration fidelity to classification performance. (3) Extensive Empirical Validation. Extensive experiments on synthetic and real-world data demonstrate that DeepFRC consistently outperforms state-of-the-art methods in both tasks, with ablation studies confirming the synergy of its components and further analyses highlighting its robustness and computational efficiency.
@@ 2 The Model
We study supervised learning with functional data, where each sample consists of an observed trajectory and a categorical label. Formally, the dataset is $\{(x_{i}(\bm{t}_{i}),y_{i})\}_{i=1}^{N}$, where $x_{i}(\bm{t}_{i})=(x_{i}(t_{i1}),\ldots,x_{i}(t_{in}))^{\intercal}$ are observations of an underlying trajectory $x(t)$ sampled at irregular and potentially misaligned time points $\bm{t}_{i}=(t_{i1},\ldots,t_{in})$, and $y_{i}\in\{1,\dots,C\}$ is its class label. We view functional data as trajectories embedded in an infinite-dimensional space, where misalignment corresponds to latent time reparameterizations. Instead of treating alignment and prediction as two separate stages, we propose to jointly learn a diffeomorphic time-warping map $\gamma:t\mapsto\tilde{t}$ and a classifier $f:x(\tilde{t})\mapsto y$ in a single end-to-end architecture (Figure 1). This allows us to directly learn representations that are invariant to temporal misalignment while remaining discriminative for classification.
Figure 1: Overview of DeepFRC. Multiple raw functional trajectories are first aligned via a neural deformation operator (1D CNN) that learns diffeomorphic time warping $\gamma(t)$, producing both warped curves and alignment functions. The aligned signals are then expanded in a Fourier basis to obtain spectral coefficients $c_{1},\dots,c_{K}$, which serve as inputs to a classifier network (MLP with Softmax output) for predicting class labels $y\in{1,\dots,C}$. Training is performed jointly by minimizing a contrastive geometric alignment loss and a classification loss.
2.1 Neural Deformation Operator for Time Warping
We introduce a neural deformation operator, parameterized by a 1D convolutional network (Kiranyaz et al., 2016; Ince et al., 2016; Kiranyaz et al., 2021), that learns velocity fields defining diffeomorphic warpings. This perspective differs from handcrafted registration methods: instead of explicitly constructing alignment rules, we parameterize $\gamma$ through a deep network optimized jointly with downstream prediction.
Given an input sequence $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n}$, a multi-layer 1D CNN extracts temporal features $\tau(x_{i}(\bm{t}_{i}))\in\mathbb{R}^{n}$:
$\bm{h}^{(l)}=\mathrm{ReLU}(W^{(l)}\circledast\bm{h}^{(l-1)}+B^{(l)}),\quad l=1,\ldots,l_{1}-1,$ (1)
with $\bm{h}^{(0)}=x_{i}(\bm{t}_{i})$. The final layer is fully connected: $\tau(x_{i}(\bm{t}_{i}))=\mathrm{ReLU}(W^{(l_{1})}\bm{h}^{(l_{1}-1)}+\bm{b}^{(l_{1})}).$ Following Chen and Srivastava (2021), we use $l_{1}=4$ convolutional blocks with kernel size $3$ and channels $16\!\rightarrow\!32\!\rightarrow\!64$. This encoder, parameterized by $\Theta_{1}$, produces latent features that encode temporal deformation fields.
To construct a boundary-preserving diffeomorphism, we transform the latent features into a monotone cumulative sum: $\tilde{\gamma}_{i}(t_{ij})=\frac{\sum_{\mu=0}^{j}\tau^{2}_{i\mu}}{\sum_{\nu=0}^{n}\tau^{2}_{i\nu}},\quad j=1,\dots,n,$ where $t_{i0}=0,t_{in}=1$, and $\tau_{i0}=0$. This ensures $\tilde{\gamma}_{i}(0)=0,\tilde{\gamma}_{i}(1)=1$, and $\frac{d\tilde{\gamma}_{i}}{dt}>0$, satisfying diffeomorphic constraints (Chen and Srivastava, 2021). To further improve smoothness and geometric consistency, we apply an additional normalization:
@@ 2.1 Neural Deformation Operator for Time Warping
This defines a neural operator that outputs valid, smooth, and data-adaptive warping functions.
2.2 Spectral Representation of Aligned Functions
Once trajectories are aligned by $\gamma_{i}$, we interpolate the aligned function $\tilde{x}_{i}(t)$ on points $\{(\gamma_{i}(t_{ij}),x_{i}(t_{ij}))\}_{j=1}^{n}$ using stable 1D linear interpolation:
$\begin{split}\tilde{x}_{i}(t)\triangleq\sum_{r=0}^{n-1}\Big[x(t_{ir})+\frac{x_{i}(t_{i(r+1)})-x_{i}(t_{ir})}{\gamma_{i}(t_{i(r+1)})-\gamma_{i}(t_{ir})}(t-\gamma_{i}(t_{ir}))\Big]\cdot\mathds{1}_{\{\gamma_{i}(t_{ir})\leq t\leq\gamma_{i}(t_{i(r+1)})\}}.\end{split}$ (3)
Rather than vectorizing $\tilde{x}_{i}(t)$ into a high-dimensional grid, which ignores smoothness and leads to inefficiency, we embed aligned functions into a compact spectral basis:
@@ 2.2 Spectral Representation of Aligned Functions
where $\{\phi_{j}(t)\}_{j=1}^{K}$ are Fourier basis functions and $c_{ij}$ are coefficients estimated by least squares:
$\tilde{\bm{c}}_{i}=\arg\!\min_{\{c_{ij}\}}\int_{0}^{1}\Big[\tilde{x}_{i}(t)-\sum_{j=1}^{K}c_{ij}\phi_{j}(t)\Big]^{2}\,dt=G^{-1}\bm{d}_{i},$ (5)
with $G_{ij}=\int_{0}^{1}\phi_{i}(t)\phi_{j}(t)\,dt$ and $d_{ij}=\int_{0}^{1}\tilde{x}_{i}(t)\phi_{j}(t)\,dt$. This spectral representation, inspired by classical functional data analysis, acts as a form of neural Fourier features, yielding compact, smooth, and regularized embeddings of aligned functions. We employ 1D linear interpolation (Eq. (3)) for its numerical stability and computational efficiency, and a Fourier basis expansion (Eq. (4)) as it introduces no additional regularization hyperparameters. Crucially, both methods satisfy the Lipschitz continuity condition required by Theorem 3.3.
2.3 Classifier Network
The spectral coefficients $\tilde{\bm{c}}_{i}\in\mathbb{R}^{K}$ are fed into a fully connected classifier with $l_{2}-1$ hidden layers and ReLU activations:
$\bm{h}^{(l)}=\mathrm{ReLU}(W^{(l)}\bm{h}^{(l-1)}+\bm{b}^{(l)}),\quad l=l_{1}+1,\ldots,l_{1}+l_{2}-1,$ (6)
followed by a softmax layer: $\psi(\tilde{\bm{c}}_{i})=\mathrm{Softmax}(W^{(l_{1}+l_{2})}\bm{h}^{(l_{1}+l_{2}-1)}+\bm{b}^{(l_{1}+l_{2})}).$ We use a three-layer MLP ($l_{2}=3$) with hidden units of size $8$ and $4$, parameterized by $\Theta_{2}$.
2.4 Objective: Contrastive Geometric Alignment and Classification
We optimize the model with parameters $\Theta=\{\Theta_{1},\Theta_{2}\}$ by combining geometric alignment and discriminative classification.
Class-aware Contrastive Alignment.
To align functional trajectories while preserving class structure, we adopt the square-root velocity function (SRVF) representation $q(t)=\text{sign}(\dot{x}(t))\sqrt{|\dot{x}(t)|}$ (Srivastava and Klassen, 2016). For a warped trajectory $x(\gamma(t))$, the SRVF is $(q\star\gamma)(t)=q(\gamma(t))\sqrt{\dot{\gamma}(t)}$. For sample $i$, the observed SRVF vector is $Q_{i}(\bm{\gamma}_{i})=((q_{i}\star\gamma_{i})(t_{i1}),\ldots,(q_{i}\star\gamma_{i})(t_{in}))^{\intercal}$. We define a contrastive alignment loss:
$\mathcal{L}_{1}(\Theta_{1})=\sum_{j=1}^{C}\frac{\sum_{i:y_{i}=j}\|Q_{i}(\bm{\gamma}_{i})-\bar{Q}^{(j)}\|}{N^{(j)}}+\alpha\sum_{1\leq u<v\leq C}\|\bar{Q}^{(u)}-\bar{Q}^{(v)}\|^{-1},$ (7)
where $\bar{Q}^{(j)}$ is the class-wise SRVF mean. The first term encourages intra-class alignment; the second is a contrastive separation term that increases inter-class margins. Together, this yields a new contrastive–geometric alignment objective.
@@ 2.4 Objective: Contrastive Geometric Alignment and Classification
The classifier is trained using standard cross-entropy:
$\mathcal{L}_{2}(\Theta)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{C}y_{ij}\log\psi_{ij}.$ (8)
Joint Objective.
The full objective integrates both alignment and prediction:
@@
where $\beta$ balances geometric alignment and classification accuracy.
Summary.
Integrating Sections 2.12.4, we obtain DeepFRC (Deep Functional Registration and Classification), an end-to-end framework coupling (i) a neural deformation operator for alignment, (ii) a spectral embedding for smooth functional representation, and (iii) a classifier guided by a contrastive–geometric loss (Figure 1). To the best of our knowledge, DeepFRC is the first model to unify alignment and prediction of functional data through diffeomorphic neural operators and contrastive geometry, offering a new paradigm for learning invariant representations of misaligned functional trajectories. A natural extension of our framework allows it to handle $d$-dimensional functional inputs ($d\geq 2$) by simply considering $x_{i}(\bm{t}_{i})\in\mathbb{R}^{n\times d}$, under the assumption of a shared warping process across dimensions, without requiring additional structural assumptions.
@@ 3 Theoretical Analysis
[…]
@@
Theorem 3.1 provides a theoretical foundation for our approach, establishing that neural networks can approximate smooth warping functions in SRVF space, thereby justifying the use of a learnable diffeomorphic registration module. While direct empirical verification of the approximation error is challenging due to the inability to obtain globally optimal warpings $\bm{\gamma}^{*}$ in real data, we validate the practical observability of this result through controlled simulation studies where the ground-truth warpings are known.
@@
[…]
@@
The assumptions of Theorem 3.3 are empirically satisfied and guide model design. Class separation $|\bar{Q}^{(u)}-\bar{Q}^{(v)}|\geq\epsilon_{0}$ is supported by distinct SRVF means (Figure 3, Table 1). The probability floor $\psi_{ij}\geq\epsilon_{0}$ is enforced via additively smoothed softmax, keeping all probabilities above $10^{-4}$. Although non-constructive, the bound informs hyperparameter tuning: $\alpha$ is chosen to maximize separation, while $\beta$ balances classification and alignment.